QuestiuII 1J1 Pom X Use the following information to determine the activation energy for the reaction shown here: A+B C +D
Answer #1

We have en (B) = E (t-1) Glven, T,= 45c= 243 +45 = 3185 T2 = 98°= 273+98= 3915 Ki = 0.0323 11 mol. s. K2 0.148 i mols Ea=7 R=thank you vmuch....